OpenStudy (anonymous):

I need some quick help with this. I took notes on my class but I don't recall being taught how to do this. Could you please help me? Problem is below: P(A) = ?, P(B) = 0.9, P(A and B) = 0.45

OpenStudy (anonymous):

I was thinking I need to set up some type of proportion

OpenStudy (phi):

is there any more info or background. It seems to be missing some info.

OpenStudy (anonymous):

Well, it says that it is an independent event.

OpenStudy (anonymous):

Events A and B are independent. Find the indicated probability. Show your work! a. P(A) = 0.3, P(B) = 0.5, P(A and B) = ? b. P(A) = ?, P(B) = 0.9, P(A and B) = 0.45

OpenStudy (phi):

in that case, (for independent events) P(A & B) = P(A)*P(B)

OpenStudy (anonymous):

That's what I did for A, but for B, the value is missing

OpenStudy (phi):

P(A) = ?, P(B) = 0.9, P(A and B) = 0.45 fill in what you know into P(A & B) = P(A)*P(B)

OpenStudy (anonymous):

For A, I got 3/20, like this: 0.3 = 3/10 0.5 = ½ 3/10 x ½ = 3/20

OpenStudy (phi):

yes for A you should get 0.15 (usually people write props as decimals)

OpenStudy (anonymous):

Oh! Okay, I'll be sure to include the decimal too.

OpenStudy (anonymous):

.45=.9 * b -->is this right?

OpenStudy (phi):

For this one P(A) = ?, P(B) = 0.9, P(A and B) = 0.45 fill in what you know into P(A & B) = P(A)*P(B) you get 0.45 = P(A)*0.9 to find P(A), divide both sides by 0.9

OpenStudy (phi):

you get \[ \frac{0.45}{0.9}= P(A)\cdot \frac{0.9}{0.9}\] which simplifies

OpenStudy (anonymous):

.45=P(A) * 9 = .5

OpenStudy (phi):

after dividing both sides by 0.9 \[ \frac{0.45}{0.9}= P(A)\cdot \frac{0.9}{0.9}\] on the right side 0.9/0.9 is 1 (anything divided by itself is one) and 1 times P(A) is P(A) (anything times 1 is itself) so you get 0.5= P(A) or P(A)= 0.5 (people usually write it this way)

OpenStudy (anonymous):

Okay so what I did is right. I got 0.5 for P(A)

OpenStudy (phi):

yes, you got it right. But notice how to do it step by step. that is the lesson to try to learn

OpenStudy (anonymous):

Thank you so much for your help. I understand. I typed every step on my paper. I'm doing a project, so I'm just going over the ones I truly couldn't understand very well. I only have two more. The rest I did myself. Again, thank you so much for your help.